ayBridget Kanen is getting ready to open a small restaurant She is on a tight budget and must choose between the foliowing long-distance phone plans LPlan A: Pay 7 cents per minute of long-distance calling Dat Plan B: Pay a fixed monthly tee of S17 for up to 260 long-distance minutes and 5 cents per minute thereafter (if she uses fewer than 260 minutes in any month, she still pays $17 for the month) Plan C: Pay a fixed monthly fee of $24 for up to 440 long-distance minutes and 3 cents per minute thereafter (f she uses fewer than 440 minutes, she still pays $24 for the month) car Read the requirements Requirement 1. Match the graphs to the total monthly costs of the three plans for dierent levels of monthly long-distance calling s d Before we match the graphs to the plans, complete the following table calculating the total costs for each plan at each of the diferent minute levels. (Round to the nearest cent.) Dat Plan 100 minutes 260 minutes 460 minutes Requirements 1. Match the graphs to the total monthly costs ot the three plans for different levels of monthly long distance calling 2. Which plan should Kanen choose if she expects to make 100 minutes of long-distance calls? 260 minutes? 460 minutes? ell T Print Done Enter any number in the edit fields and then click Check Answer.
Answer #1
Plan 100 minutes 260 minutes 460 minutes
A $             7.00 $        18.20 $        32.20
B $           17.00 $        17.00 $        27.00
C $           24.00 $        24.00 $        24.60
Best Plan A B C
Plan 100 minutes 260 minutes 460 minutes
A =100*0.07 =260*0.07 =460*0.07
B 17 17 =17+(460-260)*0.05
C 24 24 =24+(460-440)*0.03